# On geometrical origin of Kodama vector

### v2

@@ Acknowledgements.
Appendix A Curvature tensors on warped product spacetimes
In this appendix, we summarize useful relations in terms of the curvature tensor on the $D$-dimensional warped product spacetimes, ×$\mathcal{B}\times_{r}\mathcal{F}$, described by the metric (15).
@@ Appendix A Curvature tensors on warped product spacetimes
The contraction of $n$ powers of the Riemann tensor with the generalized Kronecker delta on ($D-2$)-space is given by
@@
Appendix B Formulae for curvature polynomials
In this appendix, we summarize basic properties of curvature polynomials in $D$-dimensions. (See, for example, Refs. [26, 34, 35])
@@ Appendix B Formulae for curvature polynomials
Its indices have the same properties as those of the Riemann tensor:
@@
and, in addition, it is divergence free for each index:

### v3

@@ Acknowledgements.
Appendix A Curvature tensors on warped-product spacetimes
In this appendix, we summarize useful relations in terms of the curvature tensor on the $D$-dimensional warped-product spacetimes, ×$\mathcal{B}\times_{r}\mathcal{F}$, described by the metric (15).
@@ Appendix A Curvature tensors on warped-product spacetimes
The contraction of $n$ powers of the Riemann tensor with the generalized Kronecker $\delta$ on ($D-2$)-space is given by
@@
Appendix B Formulas for curvature polynomials
In this appendix, we summarize basic properties of curvature polynomials in $D$-dimensions (see, for example, Refs. [27, 34, 35]).
@@ Appendix B Formulas for curvature polynomials
Its indices have the same properties as those of the Riemann tensor,
@@
and, in addition, it is divergence-free for each index,